# Arithmetic of cuts in ordered abelian groups and of ideals over valuation rings

### v2

@@ 2.7. The equation $S_{1}+T=S_{2}$ in ordered abelian groups of arbitrary rank
Assume that $S^{\prime}_{2}/{\mathcal{G}}_{1}=(\alpha/{\mathcal{G}}_{1})^{+}$ is principal; it then follows that $G/{\mathcal{G}}_{1}$ is discretely ordered. Hence by part 5) of Proposition 2.26, also $S_{1}/{\mathcal{G}}_{1}$ is principal. This shows that condition (*) of Theorem 2.35 holds with $S_{2}$ replaced by $S_{2}^{\prime}\,$. Now it follows from Theorem 2.35 that $T=S_{2}^{\prime}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T=S_{2}^{\prime}$. Since $S_{2}^{\prime}\subset S_{2}$, we have $S_{2}^{\prime}-_{\rm fs}S_{1}\subseteq S_{2}-_{\rm fs}S_{1}\,$. On the other hand, ${\mathcal{G}}_{1}\subseteq{\mathcal{G}}(S_{1}+(S_{2}-_{\rm fs}S_{1}))$ which implies that $S_{1}+(S_{2}-_{\rm fs}S_{1})\subseteq S_{2}^{\prime}$ and therefore $S_{2}-_{\rm fs}S_{1}\subseteq S_{2}^{\prime}-_{\rm fs}S_{1}\,$.
2): As in the proof of part 1) it is shown that $(\alpha/{\mathcal{G}}_{1})^{+}$ cannot be principal since $S_{1}/{\mathcal{G}}_{1}$ is assumed to be nonprincipal. Hence its preimage $S^{\prime}_{2}:=(\alpha+{\mathcal{G}}_{1})^{+}$ is the largest final segment contained in $S_{2}$ such that $S^{\prime}_{2}/{\mathcal{G}}_{1}$ is nonprincipal. Since it does not contain $\alpha$, it is a proper subset of $S_{2}\,$.
Again, condition (*) of Theorem 2.35 holds with $S_{2}$ replaced by $S_{2}^{\prime}$, and as before it is shown that $T=S_{2}^{\prime}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T=S_{2}^{\prime}$ and that $S_{2}^{\prime}-_{\rm fs}S_{1}\subseteq S_{2}-_{\rm fs}S_{1}\,$. To show the reverse inclusion, suppose that $S_{2}^{\prime}\subsetneq S_{1}+(S_{2}-_{\rm fs}S_{1})$. Then $(S_{1}+(S_{2}-_{\rm fs}S_{1}))/{\mathcal{G}}_{1}$ contains $\alpha/{\mathcal{G}}_{1}$ and thus $S_{1}+(S_{2}-_{\rm fs}S_{1})$ must be equal to $S_{2}\,$, which is impossible by Theorem 2.35. Hence $S_{1}+(S_{2}-_{\rm fs}S_{1})\subseteq S_{2}^{\prime}$ and therefore $S_{2}-_{\rm fs}S_{1}\subseteq S_{2}^{\prime}-_{\rm fs}S_{1}\,$. ∎
@@ 3. $\mathcal{O}$-ideals
1):  Equation (27) follows via Proposition 3.1 from the definitions of $I_{2}:I_{1}$ and $vI_{2}\,-_{\rm fs}\,vI_{1}\,$. Equation (28) follows from equation (27) since $\alpha+vI_{1}\geq 0\Leftrightarrow\alpha\geq-vI_{1}\,$. Equation (29) follows from equation (28) and the definition of $vI_{2}-_{\rm mc}vI_{1}\,$.
2):  If $I_{1}=a\mathcal{O}_{v}\,$, then $vI_{1}=(va)^{-}$ and $\Delta vI_{1}=(-va)^{-}$ by Lemma 2.8, hence by equation (29), $vI_{2}(\mathcal{O}_{v}:I_{1})=vI_{2}+\Delta vI_{1}=vI_{2}-va$. If $I_{1}$ is nonprincipal, then so is $vI_{1}$ and again by Lemma 2.8, $\Delta vI_{1}=-(vI_{1})^{c}$, whence $vI_{2}(\mathcal{O}_{v}:I_{1})=vI_{2}-(vI_{1})^{c}$ by equation (29). ∎
@@ 3.2. The invariance valuation ring of an $\mathcal{O}_{v}$-ideal
(38) $\mathcal{O}(I)\>:=\>\{b\in K\mid bI\subseteq I\}\;\;\;\mbox{\ \ and\ \ }\;\;\;\mathcal{M}(I)\>=\>\{b\in K\mid bI\subsetneq I\}\>,$
as well as
(39) $\mathcal{O}(\{0\})\>:=\>K\;\;\;\mbox{\ \ and\ \ }\;\;\;\mathcal{M}(\{0\})\>=\>\{0\}\>,$
@@
$H(I)\>:=\>H(\mathcal{O}(I))\>=\>v\mathcal{O}(I)\cap-v\mathcal{O}(I)\>.$
@@
2) For every $\mathcal{O}_{v}$-ideal $I$, $\mathcal{O}(I)^{\times}=\{b\in K\mid bI=I\}$ and $H(I)=v\mathcal{O}(I)^{\times}=H(\mathcal{O}(I))$ is a convex subgroup of $vK$.
@@
3):  By part 6) of Lemma 2.14, we compute $\mathcal{O}(I)=\{b\in K\mid bI\subseteq I\}=\{b\in K\mid vb+vI\subseteq vI\}=\{b\in K\mid vb\in{\mathcal{G}}(vI)^{-}\}=\mathcal{O}({\mathcal{G}}(vI))$, and similarly, $\mathcal{M}(I)=\{b\in K\mid bI\subsetneq I\}=\{b\in K\mid vb+vI\subsetneq vI\}=\{b\in K\mid vb\in{\mathcal{G}}(vI)^{+}\}=\mathcal{M}({\mathcal{G}}(vI))$. From what we have shown together with part 2) of our theorem, $H(I)=v\mathcal{O}(I)^{\times}=v\{b\in K\mid bI=I\}={\mathcal{G}}(vI)^{-}\setminus{\mathcal{G}}(vI)^{+}={\mathcal{G}}(vI)$.
4):  Since $\mathcal{M}(I)=\mathcal{M}({\mathcal{G}}(vI))$, this follows from part 3) and equation (36). ∎
@@
3): We have ${\mathcal{G}}(vaI)={\mathcal{G}}(va+vI)={\mathcal{G}}(vI)$ by part 3) of Lemma 2.14. Our statement follows from this by means of (41) and Proposition 3.4. ∎
@@
1): We compute: $vI\mathcal{O}=vI+v\mathcal{O}=vI+(H(\mathcal{O})\cup H(\mathcal{O})^{+})=(vI+H(\mathcal{O}))\cup(vI+H(\mathcal{O})^{+})$, where we have used Lemma 3.5. Since $H(\mathcal{O})^{+}$ consists only of positive elements, we have $vI+H(\mathcal{O})^{+}\subseteq vI$. Since $0\in H(\mathcal{O})$, we have $vI\subseteq vI+H(\mathcal{O})$. Thus, $(vI+H(\mathcal{O})^{+})\cup(vI+H(\mathcal{O}))=vI+H(\mathcal{O})$.
@@
3): By Lemma 2.25, the preimage of $wI$ in $vK$ is $vI+H(\mathcal{O})$. By part 1) of our lemma, this is equal to $vI\mathcal{O}$. ∎
@@ 3.3. Properties of $\mathcal{M}(I)$
7): Since $v(I:\mathcal{M}(I))=\{va\mid va+v\mathcal{M}(I)^{+}\subseteq vI\}$ by (27) of Lemma 3.3 and $v\mathcal{M}(I)={\mathcal{G}}(vI)^{+}$, the statement follows from part 2) of Proposition 2.20.
8): We have that $I\mathcal{M}(I)\subsetneq I$ if and only if $vI+{\mathcal{G}}(vI)^{+}\neq vI$. By Lemma 2.27, this holds if and only if $vI/{\mathcal{G}}(vI)$ has a smallest element. Denote the valuation associated with $\mathcal{O}$ by $w$. Since ${\mathcal{G}}(vI)=H(\mathcal{O}(I))$ by (42), this in turn holds if and only if $wI=vI/H(\mathcal{O}(I))=vI/{\mathcal{G}}(vI)$ has a smallest element, which means that $I$ is a principal $\mathcal{O}(I)$-ideal.
9): The assumption $\mathcal{M}(I)\subsetneq\mathcal{M}(J)$ is equivalent to $\mathcal{O}(J)\subsetneq\mathcal{O}(I)$. By (41) and Proposition 3.4, this is equivalent to ${\mathcal{G}}(vJ)\subsetneq{\mathcal{G}}(vI)$. By part 3) of Proposition 2.20, there is $\alpha\in vK$ such that $vI+vJ=\alpha+vI$. Choosing $a\in K$ with $va=\alpha$, we obtain that $IJ=aI$.
@@
12): By (29), $vI(\mathcal{O}_{v}:I)=vI-_{\rm mc}vI$. If $vI$ is nonprincipal, then by part 1) of Corollary 2.19, $vI-_{\rm mc}vI={\mathcal{G}}(vI)^{+}$, so $I(\mathcal{O}_{v}:I)=\mathcal{M}(I)$. If $vI$ is principal, then by Lemma 2.12, $vI-_{\rm mc}vI=0^{-}$, so $I(\mathcal{O}_{v}:I)=\mathcal{O}_{v}\,$.
@@ 3.4. The equation $I_{1}J=I_{2}$
Define $w$ as above. Applying (28) of Lemma 3.3 twice, we obtain $w(\mathcal{O}:(\mathcal{O}:J))=\Delta w(\mathcal{O}:J)=\Delta(\Delta wJ)$. By part 6) of Lemma 2.10, this is equal to the closure of $wJ$ in $wK$. Our assertion now follows by Proposition 3.1. ∎
@@
1): By Theorem 3.6, $J$ is an $\mathcal{O}(J)$-ideal. Denote by $w$ the valuation associated with $\mathcal{O}(J)$, so that $wa=va/H(\mathcal{O}(J))$ for every $a\in K$. By (42), $H(\mathcal{O}(J))={\mathcal{G}}(vJ)$. By definition, $(vJ)^{\diamondsuit}$ is the preimage of $\widehat{vJ/{\mathcal{G}}(vJ)}=\widehat{vJ/H(\mathcal{O}(J))}=\widehat{wJ}$ in $vK$. If $wJ$ is closed in $wK$, then $\widehat{wJ}=wJ=vJ/{\mathcal{G}}(vJ)$ whose preimage is $vJ+{\mathcal{G}}(vJ)=vJ$ and therefore, $I_{(vJ)^{\diamondsuit}}=I_{vJ}=J=J^{\diamondsuit}$. Now assume that $wJ$ is not closed in $wK$ and take $a\in K$ such that $wa$ is the infimum of $wJ$ in $wK$. Then $\widehat{wJ}=(wa)^{-}$ whose preimage is $va^{-}+{\mathcal{G}}(vJ)$. By part 2) of Lemma 3.8, this is equal to $va\mathcal{O}(I)$. It follows that $I_{(vJ)^{\diamondsuit}}=I_{vaO(I)}=aO(I)=J^{\diamondsuit}$. This proves our assertion.
@@
()(*)( ∗ )  $\mathcal{O}(I_{1})\subseteq\mathcal{O}(I_{2})$, and $I_{2}$ is a nonprincipal $\mathcal{O}(I_{1})$-ideal if $I_{1}$ is a nonprincipal $\mathcal{O}(I_{1})$-ideal.
holds, then $J=I_{2}-I_{1}$ is the largest solution of equation (6), and it is equal to:
@@
ii) $\mathcal{O}(I_{1})=\mathcal{O}(I_{2})$ and $I_{1}$ is a nonprincipal $\mathcal{O}(I_{1})$-ideal, but $I_{2}$ is a principal $\mathcal{O}(I_{1})$-ideal
@@
Since $\mathcal{O}(I)=\mathcal{O}({\mathcal{G}}(vI))$ for every $\mathcal{O}_{v}$-ideal $I$ by equation (41), we know that $\mathcal{O}(I_{1})\subseteq\mathcal{O}(I_{2})$ if and only if ${\mathcal{G}}(vI_{1})\subseteq{\mathcal{G}}(vI_{2})$. Part 2) of Lemma 3.2, with the valuation $w$ associated with $\mathcal{O}(I_{1})$ in place of $v$, shows for $i=1,2$ that $I_{i}$ is a principal $\mathcal{O}(I_{1})$-ideal if and only if $wI_{i}=vI_{i}/{\mathcal{G}}(I_{1})$ is principal. Further, if $vI_{1}/{\mathcal{G}}(I_{1})$ is nonprincipal, then also $I_{1}$ is nonprincipal and therefore, $vI_{2}(\mathcal{O}_{v}:I_{1})=vI_{2}-(vI_{1})^{c}$ by (30).
@@
In both cases, $J=I^{\prime}_{2}:I_{1}$ is the largest $\mathcal{O}_{v}$-ideal such that $I_{1}\,J\subseteq I_{2}\,$; it is the largest solution of $I_{1}\,J=I^{\prime}_{2}\,$.

### v3

@@ 2.7. The equation $S_{1}+T=S_{2}$ in ordered abelian groups of arbitrary rank
Assume that $S^{\prime}_{2}/{\mathcal{G}}_{1}=(\alpha/{\mathcal{G}}_{1})^{+}$ is principal; it then follows that $G/{\mathcal{G}}_{1}$ is discretely ordered. Hence by part 5) of Proposition 2.25, also $S_{1}/{\mathcal{G}}_{1}$ is principal. This shows that condition ()(*)( ∗ ) of Theorem 2.33 holds with $S_{2}$ replaced by $S_{2}^{\prime}\,$. Now it follows from Theorem 2.33 that $T=S_{2}^{\prime}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T=S_{2}^{\prime}$. Since $S_{2}^{\prime}\subset S_{2}$, we have $S_{2}^{\prime}-_{\rm fs}S_{1}\subseteq S_{2}-_{\rm fs}S_{1}\,$. On the other hand, ${\mathcal{G}}_{1}\subseteq{\mathcal{G}}(S_{1}+(S_{2}-_{\rm fs}S_{1}))$ which implies that $S_{1}+(S_{2}-_{\rm fs}S_{1})\subseteq S_{2}^{\prime}$ and therefore $S_{2}-_{\rm fs}S_{1}\subseteq S_{2}^{\prime}-_{\rm fs}S_{1}\,$.
2): As in the proof of part 1) it is shown that $S^{\prime}_{2}\subsetneq S_{2}$ and that ${\mathcal{G}}(S^{\prime}_{2})={\mathcal{G}}_{1}$. Also as in the proof of part 1) it is shown that if $S^{\prime}_{2}/{\mathcal{G}}_{1}=(\alpha/{\mathcal{G}}_{1})^{+}$ is principal, then so is $S_{1}/{\mathcal{G}}_{1}$. But this contradicts our assumption, so $(\alpha/{\mathcal{G}}_{1})^{+}$ is nonprincipal. Since $\alpha+{\mathcal{G}}_{1}=\alpha+{\mathcal{G}}(S_{2})$ is an initial segment od $S_{2}\,$, any final segment $S$ that properly contains $S^{\prime}_{2}$ will contain an element $\beta\in\alpha+{\mathcal{G}}_{1}\,$. This implies that $S/{\mathcal{G}}_{1}$ contains $S_{2}/{\mathcal{G}}_{1}\,$, so if it is contained in $S_{2}/{\mathcal{G}}_{1}\,$, then it is equal to $S_{2}/{\mathcal{G}}_{1}$ and hence principal. This proves that $S^{\prime}_{2}$ is the largest final segment contained in $S_{2}$ such that $S^{\prime}_{2}/{\mathcal{G}}_{1}$ is nonprincipal.
Again, condition ()(*)( ∗ ) of Theorem 2.33 holds with $S_{2}$ replaced by $S_{2}^{\prime}$ so that $T=S_{2}^{\prime}-_{\rm fs}S_{1}$ is the largest solution of $S_{1}+T=S_{2}^{\prime}$ and as before, $S_{2}^{\prime}-_{\rm fs}S_{1}\subseteq S_{2}-_{\rm fs}S_{1}\,$. To show the reverse inclusion, suppose that $S_{2}^{\prime}\subsetneq S_{1}+(S_{2}-_{\rm fs}S_{1})$. Then $(S_{1}+(S_{2}-_{\rm fs}S_{1}))/{\mathcal{G}}_{1}$ contains $\alpha/{\mathcal{G}}_{1}$ and thus $S_{1}+(S_{2}-_{\rm fs}S_{1})$ must be equal to $S_{2}\,$, which is impossible by Theorem 2.33. Hence again, $S_{1}+(S_{2}-_{\rm fs}S_{1})\subseteq S_{2}^{\prime}$ and therefore $S_{2}-_{\rm fs}S_{1}\subseteq S_{2}^{\prime}-_{\rm fs}S_{1}\,$. ∎
@@ 3. $\mathcal{O}$-ideals
1):  Equation (33) follows via Proposition 3.1 from the definitions of $I_{2}:I_{1}$ and $vI_{2}\,-_{\rm fs}\,vI_{1}\,$. Equation (34) follows from equation (33) since $\alpha+vI_{1}\geq 0\Leftrightarrow\alpha\geq-vI_{1}\,$. Equation (35) follows from equation (34) and part 1) of Lemma 2.10.
2):  If $I_{1}=a\mathcal{O}_{v}\,$, then $vI_{1}=(va)^{-}$ and $\Delta vI_{1}=(-va)^{-}$ by Lemma 2.8, hence by equations (35) and (15), $vI_{2}(\mathcal{O}_{v}:I_{1})=vI_{2}+\Delta vI_{1}=vI_{2}-va$. If $I_{1}$ is nonprincipal, then so is $vI_{1}$ and again by Lemma 2.8, $\Delta vI_{1}=-(vI_{1})^{c}$, whence $vI_{2}(\mathcal{O}_{v}:I_{1})=vI_{2}-(vI_{1})^{c}$ by equation (35). ∎
@@ 3.2. The invariance valuation ring of an $\mathcal{O}_{v}$-ideal
(44) $\mathcal{O}(I)\>:=\>\{b\in K\mid bI\subseteq I\}\;\;\;\mbox{\ \ and\ \ }\;\;\;\mathcal{M}(I)\>=\>\{b\in K\mid bI\subsetneq I\}\>.$
@@
$H(I)\>:=\>H(\mathcal{O}(I))\>=\>v\mathcal{O}(I)\cap-v\mathcal{O}(I)\>=\>v\mathcal{O}(I)^{\times}\>.$
@@
2) For every $\mathcal{O}_{v}$-ideal $I$, $\mathcal{O}(I)^{\times}=\{b\in K\mid bI=I\}$ and $H(I)$ is a convex subgroup of $vK$.
@@
3):  By part 6) of Lemma 2.13, we compute $\mathcal{O}(I)=\{b\in K\mid bI\subseteq I\}=\{b\in K\mid vb+vI\subseteq vI\}=\{b\in K\mid vb\in{\mathcal{G}}(vI)^{-}\}=\mathcal{O}({\mathcal{G}}(vI))$, and similarly, $\mathcal{M}(I)=\{b\in K\mid bI\subsetneq I\}=\{b\in K\mid vb+vI\subsetneq vI\}=\{b\in K\mid vb\in{\mathcal{G}}(vI)^{+}\}=\mathcal{M}({\mathcal{G}}(vI))$. From what we have shown, $H(I)=v\mathcal{O}(I)^{\times}=v\{b\in K\mid bI=I\}={\mathcal{G}}(vI)^{-}\setminus{\mathcal{G}}(vI)^{+}={\mathcal{G}}(vI)$.
4):  Since $\mathcal{M}(I)=\mathcal{M}({\mathcal{G}}(vI))$, this follows from part 3) and (42). ∎
@@
3): We have ${\mathcal{G}}(vaI)={\mathcal{G}}(va+vI)={\mathcal{G}}(vI)$ by part 3) of Lemma 2.13. Our statement follows from this by means of (46). ∎
@@
1): We compute: $vI\mathcal{O}=vI+v\mathcal{O}=vI+(H(\mathcal{O})\cup H(\mathcal{O})^{+})=(vI+H(\mathcal{O}))\cup(vI+H(\mathcal{O})^{+})$, where we have used Lemma 3.5. Since $H(\mathcal{O})^{+}$ consists only of positive elements, we have $vI+H(\mathcal{O})^{+}\subseteq vI$. Since $0\in H(\mathcal{O})$, we have $vI\subseteq vI+H(\mathcal{O})$. Thus, $(vI+H(\mathcal{O}))\cup(vI+H(\mathcal{O})^{+})=vI+H(\mathcal{O})$.
@@
3): By (37) and Lemma 2.24, the preimage of $wI$ under $\varphi_{H(\mathcal{O})}$ is $vI+H(\mathcal{O})$. By part 1) of our lemma, this is equal to $vI\mathcal{O}$. ∎
@@ 3.3. Properties of $\mathcal{M}(I)$
7): Since $v(I:\mathcal{M}(I))=\{va\mid va+v\mathcal{M}(I)\subseteq vI\}$ by (33) of Lemma 3.3 and $v\mathcal{M}(I)={\mathcal{G}}(vI)^{+}$, the statement follows from part 2) of Proposition 2.20.
8): We have that $I\mathcal{M}(I)\subsetneq I$ if and only if $vI+{\mathcal{G}}(vI)^{+}\neq vI$. By Lemma 2.26, this holds if and only if $vI/{\mathcal{G}}(vI)$ has a smallest element. Denote the valuation associated with $\mathcal{O}(I)$ by $w$. Since ${\mathcal{G}}(vI)=H(\mathcal{O}(I))$ by (47), this in turn holds if and only if $wI=vI/H(\mathcal{O}(I))=vI/{\mathcal{G}}(vI)$ has a smallest element, which means that $I$ is a principal $\mathcal{O}(I)$-ideal.
9): The assumption $\mathcal{M}(I)\subsetneq\mathcal{M}(J)$ is equivalent to $\mathcal{O}(J)\subsetneq\mathcal{O}(I)$. By (46) and Proposition 3.4, this is equivalent to ${\mathcal{G}}(vJ)\subsetneq{\mathcal{G}}(vI)$. By part 3) of Proposition 2.20, there is $\alpha\in vK$ such that $vI+vJ=vI+\alpha$. Choosing $a\in K$ with $va=\alpha$, we obtain that $IJ=aI$.
@@
12): By (35), $vI(\mathcal{O}_{v}:I)=vI-_{\rm mc}vI$. If $vI$ is nonprincipal, then by part 1) of Corollary 2.19, $vI-_{\rm mc}vI={\mathcal{G}}(vI)^{+}$, so $I(\mathcal{O}_{v}:I)=\mathcal{M}(I)$. If $vI$ is principal, then from (19) with $\alpha=\beta$ we obtain $vI-_{\rm mc}vI=0^{-}$, so $I(\mathcal{O}_{v}:I)=\mathcal{O}_{v}\,$.
@@ 3.4. The equation $I_{1}J=I_{2}$
Define $w$ as above. Applying (34) of Lemma 3.3 twice, we obtain $w(\mathcal{O}:(\mathcal{O}:J))=\Delta w(\mathcal{O}:J)=\Delta(\Delta wJ)$. By part 3) of Lemma 2.8, this is equal to the closure of $wJ$ in $wK$. Our assertion now follows by Proposition 3.1. ∎
@@
1): By Theorem 3.6, $J$ is an $\mathcal{O}(J)$-ideal. Denote by $w$ the valuation associated with $\mathcal{O}(J)$, so that $wa=va/H(\mathcal{O}(J))$ for every $a\in K$. By (47), $H(\mathcal{O}(J))={\mathcal{G}}(vJ)$. By definition, $(vJ)^{\diamondsuit}$ is the preimage of $\widehat{vJ/{\mathcal{G}}(vJ)}=\widehat{vJ/H(\mathcal{O}(J))}=\widehat{wJ}$ under $\varphi_{H(\mathcal{O}(J))}$. If $wJ$ is closed in $wK$, then $\widehat{wJ}=wJ=vJ/{\mathcal{G}}(vJ)$ whose preimage is $vJ+{\mathcal{G}}(vJ)=vJ$ and therefore, $I_{(vJ)^{\diamondsuit}}=I_{vJ}=J=J^{\diamondsuit}$. Now assume that $wJ$ is not closed in $wK$ and take $a\in K$ such that $wa$ is the infimum of $wJ$ in $wK$. Then $\widehat{wJ}=(wa)^{-}$ whose preimage is $va^{-}+{\mathcal{G}}(vJ)$. By part 2) of Lemma 3.8, this is equal to $va\mathcal{O}(I)$. It follows that $I_{(vJ)^{\diamondsuit}}=I_{vaO(I)}=aO(I)=J^{\diamondsuit}$. This proves our assertion.
@@
()(*)( ∗ )  $\mathcal{O}(I_{1})\subseteq\mathcal{O}(I_{2})$, and $I_{2}$ is a nonprincipal $\mathcal{O}(I_{2})$-ideal if $I_{1}$ is a nonprincipal $\mathcal{O}(I_{2})$-ideal
holds, then $J=I_{2}:I_{1}$ is the largest solution of equation (6), and it is equal to:
@@
ii) $\mathcal{O}(I_{1})=\mathcal{O}(I_{2})$ and $I_{1}$ is a nonprincipal $\mathcal{O}(I_{1})$-ideal, but $I_{2}$ is a principal
$\ \mathcal{O}(I_{1})$-ideal
@@
Since $\mathcal{O}(I)=\mathcal{O}({\mathcal{G}}(vI))$ for every $\mathcal{O}_{v}$-ideal $I$ by equation (46), we know that $\mathcal{O}(I_{1})\subseteq\mathcal{O}(I_{2})$ if and only if ${\mathcal{G}}(vI_{1})\subseteq{\mathcal{G}}(vI_{2})$. Part 2) of Lemma 3.2, with the valuation $w$ associated with $\mathcal{O}(I_{2})$ in place of $v$, shows for $i=1,2$ that $I_{i}$ is a principal $\mathcal{O}(I_{2})$-ideal if and only if $wI_{i}=vI_{i}/{\mathcal{G}}(I_{2})$ is principal. Further, if $vI_{1}/{\mathcal{G}}(I_{2})$ is nonprincipal, then also $I_{1}$ is nonprincipal and therefore, $vI_{2}(\mathcal{O}_{v}:I_{1})=vI_{2}-(vI_{1})^{c}$ by (36).
@@
In both cases, $J=I^{\prime}_{2}:I_{1}$ is the largest solution of $I_{1}\,J=I^{\prime}_{2}\,$, and $I^{\prime}_{2}:I_{1}=I_{2}:I_{1}\,$.